A sample of gas was found to occupy 37.82 L under conditions of 432 kPa and 453 K.  What is the identity of this diatomic gas? The mass of the gas was found to be 139 grams.


1
Expert's answer
2020-10-26T14:49:21-0400

sample of gas was found to occupy 37.82 L


under conditions of 432 kPa and 453 K. 


The mass of the gas was found to be 139 grams.


Therefore:


((37.82)(432))/ 453 = 139



16338.24 b /453 = 139


36.07b = 139


b = 139/36.07


b = 3.85g


MCo2 = 3.85g